I am working on this question and need help.
Use the intermediate value theorem to determine whether g(x)=4x^3-3x+3 has a zero between -2 and -1
Any help you can give me is greatly appreciated.
Thanks in Advance!

Answer by oscargut(2103)

g(x)=4x^3-3x+3
g(-2)=-32+6+3=-23
g(-1)=-4+3+3=2
then g has a zero between -2 and -1 because g(-2)<0 and g(-1)>0